OpenStudy (anonymous):

if f(x)={(3,5),(2,4),(1,7)} and g(x)=square root(x-3) what do i do to determine (f+g)(1)

OpenStudy (anonymous):

well first, (f+g)(1) = f(1)+g(1), so the name of the game is to figure out what f(1) and g(1) is and add them

OpenStudy (anonymous):

but i dont understand how to find f(x) if theres nothing to solve for its only sets of numbers

OpenStudy (anonymous):

those points are solutions. For example if i have a function h(x), and i say: h(x) = {(1,2), (3,14), (5, 8)} What im really saying is: h(1) = 2, h(3) = 14, h(5) = 8

OpenStudy (anonymous):

so in this case f(1) is seven and for the other one the g(1) i just get -2

OpenStudy (anonymous):

you are correct on f(1), that is seven :) for g(1), you have: \[g(x) = \sqrt{x-3}\] in your problem above, was it supposed to be just x - 3 without the square root?

OpenStudy (anonymous):

no its including the square root ...can it be a negative 2 inside the square root

OpenStudy (anonymous):

if its negative inside the square root, you are gonna need imaginary (complex) numbers. \[g(1) = \sqrt{1-3} \Rightarrow g(1) = \sqrt{-2} \Rightarrow g(1) = i\sqrt{2}\]